# Unsupervised Concept Drift Detection based on Parallel Activations of Neural Network

### v1

@@ 4 Experimental evaluation
The experiment compared the performance of the proposed approach with reference methods. The results for streams with ten drifts are shown in Figure 2. The columns indicate the results for different numbers of features, and the rows for gradual (top) and sudden drift (bottom) types. On the horizontal axis of each graph, successive chunks of the data stream are visible, while the central moments of the actual drift are marked with ticks and grid. Each detection is marked with a single point. For emphasis, the proposed approach is shown in red. The consecutive lines show the results from subsequent replications for a given detector. General results in the form of drift detection error measures averaged over the stream replications are presented in Figure 3.
@@
Figure 2: Detection moments of all evaluated methods for streams with 10 drifts. The actual drifts are marked with ticks on horizontal axis and particular methods’s detections in ten stream replications with black or red (in case of the proposed approach) points.
@@
These graphs allow to visually assess the detection quality. The ideal result would be for the method to flag a single change at the center point of the each drift occurrence across all replications – this would result in vertical lines overlapping with the drift markings on the horizontal axis. It can be seen that detections for all methods are more dispersed in the case of gradual drifts – this is either due to the detection in an early or later phase of drift, or due to multiple signaling of the same change in the case of high sensitivity of the method.
@@
Looking at the color-coded tables in the form of heat maps (Figure 3) allows for more general interpretation. The low error values are marked in blue and high in red. Cases for which it was impossible to calculate errors – due to lack of detection – are marked as blank fields. The columns of the figure show the results for the three drift detection errors – $D1$, $D2$, and $R$, respectively. In each heat map, the results for individual methods are presented in columns, and the results for all tested types of streams are presented in rows, averaged over ten replications.
@@
Figure 3: Drift detection error measures: D1 (left), D2 (center) and R (right) for all evaluated detection methods (horizontal axis) and all evaluated data streams (vertical axis). High error values are marked in red color; low values in blue.
@@
In the first error, the average distance of each detection to the nearest drift ($D1$), the average distance to the nearest drift is measured for each detection. This means that methods that signal many distant drifts from the moment of their actual occurrence, such as ddm, eddm, and md3, will achieve higher error values. It is also visible that for streams where drifts are less frequent (3 and 5 in the entire stream), these errors’ values will be higher due to the relatively greater distance for incorrect detections. In the second measure, the average distance of each drift to the nearest detection ($D2$), the nearest detection is searched for each drift. The highest errors in this criterion are demonstrated by the ddm and md3 methods, which sometimes signaled drifts with a delay or did not detect all changes. The last measure, the adjusted ratio of the number of detections to the number of drifts ($R$), looks only at the number of changes signaled by the method compared to the actual number of drifts. This measure is aimed to penalize methods for too many detections rather than too few – hence the high error values for the md3, ddm, and eddm methods. It is worth paying attention to the results for the cddd method, which are exceptionally low in all streams for which the measure could be determined, however, its problem is the tendency to not signal frequent and gradual drifts.
@@
The results obtained from the second experiment were also subjected to post-hoc Nemenyi test based on Wilcoxon Signed Rank across all the three errors. The results for $D1$, $D2$, and $R$ errors are presented in the Figure 4. The overall statistical analysis did not consider the results of the cddd detector due to the inability to calculate measures in the absence of detection by the method, which took place in 12 out of 24 calculations.
@@
Figure 4: Critical Difference diagrams for three drift detection error measures: D1 (left), D2 (center) and R (right). The cddd method was not taken into account in the statistical evaluation due to lack of possibility to calculate drift detection measures in certain data streams.
The results show that in the case of $D1$ error, the presented method is significantly best, lone leader of comparison. For the $R$ error, the presented solution – while having the best ranking value – is statistically significantly co-dependent to the adwin and occd. For the $D2$, the best results were obtained by the ocdd method, and the padd results are statistically dependent on the second method in terms of quality, the adwin method with the minimal difference in ranks.
It is worth to emphasize, that those criteria should not be used independently to evaluate methods, and it is the juxtaposition of all three that describes the proper and effective method operation.
@@ 5 Conclusions and future works
This work proposed an unsupervised Parallel Activations Drift Detector, that utilises an untrained neural network to recognize significant changes in the posterior distribution of the data stream to signal concept drift according to the result of a statistical test stabilized by pool of replications. The conducted experimental evaluation allowed to demonstrate that the proposed padd method states a valuable tool in the context of reference methods. This enriches the pool of available drift detection methods, introducing an algorithm with high reliability in a proper configuration.
@@ Acknowledgements
This work was supported by the statutory funds of the Department of Systems and Computer Networks, Wroclaw University of Science and Technology.

### v2

@@ 4.2 Comparison with reference methods
The results for streams with ten drifts are shown in Figure 4. The columns indicate the results for different numbers of features – from 30 in the first column to 90 in the last one – and the rows for different numbers of drifts present in the stream – from 3 drifts in the first row to 15 drifts in the last row. On the horizontal axis of each graph, successive chunks of the data stream are visible, while the central moments of the actual drift are marked with ticks on the x-axis and a grid. Each detection is marked with a single point. For emphasis, the proposed approach is shown in red. The consecutive lines show the results from subsequent replications for a given detector. The last row shows the concept drift dynamics.
@@
Figure 4: Detection moments of all evaluated methods for streams with sudden concept drifts. The actual drifts are marked with ticks on the horizontal axis and particular methods’s detections in ten stream replications with black or red (in the case of the proposed approach) points. The ideal drift detection would result in vertical lines overlapping with ticks on x-axis.
@@
These graphs allow to visually assess the detection quality. The ideal result would be for the method to flag a single change at the center point of each drift occurrence across all replications – this would result in vertical lines overlapping with the drift markings on the horizontal axis.
Figure 5 shows the results for streams with gradual concept drift similarly. It can be seen that detections for all methods are more dispersed in the case of gradual drifts – this is either due to the detection in an early or later phase of drift or due to multiple signaling of the same drift stretched over time in the case of high sensitivity of the method.
Figure 5: Detection moments of all evaluated methods for streams with gradual concept drifts. The ticks on x-axis indicate the central point of each concept drift.
@@
Looking at the color-coded tables in the form of heat maps, shown in Figure 6, allows for a more general interpretation. The low error values are marked in blue and high in red. Cases for which it was impossible to calculate errors – due to lack of detection – are left as blank cells. The columns of the figure show the results for the three drift detection error measures – $D1$, $D2$, and $R$, respectively. In each heat map, the results for individual methods are presented in columns, and the results for all tested types of streams are presented in rows, averaged over ten replications.
@@
Figure 6: Drift detection error measures: D1 (left), D2 (center), and R (right) for all evaluated detection methods (horizontal axis) and all evaluated data streams (vertical axis). High error values are marked in red color; low values in blue.
@@
In the first error, the average distance of each detection to the nearest drift ($D1$), the average distance to the nearest drift is measured for each detection. This means that methods that signal many distant drifts from the moment of their actual occurrence, such as ddm, eddm, and md3, will achieve higher error values. It is also visible that for streams where drifts are less frequent (3 and 5 in the entire stream), these errors’ values will be higher due to the relatively greater distance for incorrect detections.
In the second measure, the average distance of each drift to the nearest detection ($D2$), a single nearest detection is searched for each drift. The $D2$ measure will penalize the methods that fail to recognize the actually occurring drifts or do so with a delay. The highest errors in this criterion are demonstrated by the ddm and md3 methods, which sometimes signaled drifts with a delay or did not detect all changes.
The last measure, the adjusted ratio of the number of drifts to the number of detections ($R$), looks only at the actual number of drifts compared to the number of changes signaled by the method. In the case of too many detections, the measure will return values close to (but not exceeding) one. However, when the number of signaled changes is lower than the number of drifts, the measure will return values exceeding one. High error values are visible for the ddm and eddm methods, but they do not exceed one, hence the methods were penalized for too many detections. However, the md3 method obtained measure values exceeding one, which indicates too few detections on average, compared to the number of drifts. It is worth paying attention to the results for the cddd method, which are exceptionally low in all streams for which the measure could be determined, but its problem is the tendency to not signal frequent and gradual drifts.
@@
The results obtained from the second experiment were also subjected to post-hoc Nemenyi test based on Wilcoxon Signed Rank across all three errors. The results for $D1$, $D2$, and $R$ errors are presented in the Figure 7. The overall statistical analysis did not consider the results of the cddd detector due to the inability to calculate measures in the absence of detection by the method, which took place in 12 out of 24 calculations.
@@
Figure 7: Critical Difference diagrams for three drift detection error measures: D1 (left), D2 (center), and R (right). The cddd method was not taken into account in the statistical evaluation due to the lack of possibility to calculate drift detection measures in certain data streams.
The results show that in the case of $D1$ error, the presented method is significantly the best lone leader of comparison. For the $R$ error, the presented solution – while having the best ranking value – is statistically significantly co-dependent to the adwin and occd. For the $D2$, the best results were obtained by the ocdd method, and the padd results are statistically dependent on the second method in terms of quality, the adwin method with the minimal difference in ranks.
It is worth emphasizing that those criteria should not be used independently to evaluate methods, and it is the juxtaposition of all three that describes the proper and effective method operation.
@@ 5 Conclusions and future works
This work proposed an unsupervised Parallel Activations Drift Detector that utilizes an untrained neural network to recognize significant changes in the posterior distribution of the data stream to signal concept drift according to the result of a statistical test stabilized by a pool of replications. The conducted experimental evaluation allowed to demonstrate that the proposed padd method states a valuable tool in the context of reference methods. This enriches the pool of available drift detection methods, introducing an algorithm with high reliability in a proper configuration.
@@ Acknowledgements
This work was supported by the statutory funds of the Department of Computer Systems and Networks, Wroclaw University of Science and Technology.